# Time-tronics: from temporal printed circuit board to quantum computer

### v2

@@
thanks: krzysztof.sacha@uj.edu.pl
Time-tronics: from temporal printed circuit board to quantum computer
[…]
Time crystalline structures can be created in periodically driven systems [1, 2]. They are temporal lattices which can reveal different condensed matter behaviours ranging from Anderson localization in time to temporal analogues of many-body localization or topological insulators [1, 3]. However, the potential practical applications of time crystalline structures have yet to be explored. Here, we pave the way for time-tronics where temporal lattices are like printed circuit boards for realization of a broad range of quantum devices. The elements of these devices can correspond to structures of dimensions higher than three and can be arbitrarily connected and reconfigured at any moment. Moreover, our approach allows for the construction of a quantum computer, enabling quantum gate operations for all possible pairs of qubits. Our findings indicate that the limitations faced in building devices using conventional spatial crystals can be overcome by adopting crystalline structures in time.
@@
We first illustrate how resonantly driven ultra-cold atoms can create a time crystalline structure where the connection of arbitrary sites through the tunneling of atoms between sites or atom-atom interactions can be realized. As any connections between sites can be controlled, it is possible to realize a broad class of quantum devices, ranging from one-, two-, three- or higher-dimensional structures to more exotic objects like a Klein bottle, all of which can be connected in an arbitrary way, and the entire system can be reconfigured at any time during an experiment.
Next, we focus on a specific example, a quantum computer [39]. We demonstrate that a temporal printed circuit board can host qubits, where all single-qubit operations can be realized and a controlled-Z gate can be performed between all possible qubit pairs, meeting the conditions for a universal quantum computer.
Temporal printed circuit board. To explain the creation of a time crystalline structure, let us begin with a classical particle moving periodically in a one-dimensional (1D) box between infinite walls. Such a periodic trajectory remains stable in the presence of a periodically oscillating potential if the resonance condition is met, i.e., if the frequency $\omega$ of the oscillating potential is $s$ times greater than the frequency of the particle motion, where $s$ is an integer number [29]. There are $s$ oscillations of the potential for each period of the particle trajectory, allowing for the simultaneous driving of $s$ particles that follow each other with a delay $T=2\pi/\omega$, as illustrated in Fig. 1a (see also Methods).
@@
Figure 1: (a): Red ball illustrates a classical particle moving in a 1D box potential in the presence of an oscillating spatially periodic potential. If the period of particle motion is $s$ times longer than the period of the oscillating potential, $T=2\pi/\omega$, we have an $s:1$ resonance. In such a case, $s$ classical non-interacting particles can be positioned so that they follow one another along the resonant trajectory. In the quantum description, the $s:1$ resonance manifests itself as $s$ localized wave-packets, which evolve one after another along the resonant trajectory. Here, we assume that the wave-packets are sufficiently localized so that natural tunneling between neighboring wave-packets is negligible. (b): If during the encounter of the $i$-th and $j$-th wave-packets moving in opposite directions, one broad laser beam (labelled $\mathbf{k}_{2}$) and one narrow laser beam (labelled $\mathbf{k}_{1}$) focused on the meeting point of the wave-packets are briefly turned on and the Bragg scattering [40] condition is satisfied, we can realize atom tunneling between the wave-packets with an amplitude $J_{ij}$, the modulus of which depends on the beam parameters and the interaction time with the atom, and a phase depending on the relative phase between the beams. With a controllable array of focused laser beams available (e.g., using a digital micromirror device [41] as in the figure), which can be independently activated at different time points, one can control the tunneling amplitudes $J_{ij}$ between any pair of wave-packets. (c): Ultra-cold atoms can be prepared via a Feshbach resonance in states where they do not interact. In such a situation, we have the possibility of selectively controlling interactions between atoms occupying any pair of wave-packets. If just before the encounter of atoms occupying two wave-packets, we perform a Raman transfer [42] (using broad and focused laser beams) from the initial internal states of the atoms to states where the atoms interact, then during the passage of the wave-packets, interaction between atoms will occur. After the wave-packets pass each other, we perform a Raman transfer again, but back to non-interacting internal states of the atoms. With a digital micromirror device available, interaction between atoms occupying any pair of wave-packets can be realized. (d): Error of atom transfer between two wave-packets using Bragg pulses versus the assumed number of cycles of the resonant trajectory needed to achieve full transfer (two Bragg pulses per each cycle). Faster transfer requires stronger Bragg pulses, $\lambda_{\rm Bragg}$, increasing the coupling of the atom to other undesired states. Generally, apart from additional coherent oscillations, the longer the realization of the transfer, the smaller the error. (e) and (f): Fidelity of the controlled-Z gate (CZ), in which the interaction imparts a phase of $\pi$ to the state where one atom occupies one wave-packet and the second atom occupies another wave-packet, versus the number of required cycles (two interaction meetings of the wave-packets per cycle). The two curves represent the cases without (black) and with (red) the inclusion of Raman transfer errors [43] (see Methods). For a small number of cycles, stronger interactions are needed, and the coupling to other states limits the CZ fidelity. For a larger number of cycles, more Raman transfers are needed, and their imperfections limit the fidelity. (e) is for a variable interaction $g_{0}$ and fixed Raman transfer time ($T_{\rm Raman}\approx 0.12T$, before and after the center-point meeting time) and (f) is for a fixed interaction ($g_{0}=10$) and variable Raman transfer time. The results presented in (d)-(f) correspond to 39K atoms driven resonantly by an optical lattice potential, created by laser radiation with a wavelength $10.6\;\mu$m, which oscillates with a frequency 5.46 kHz and with Bragg pulses of wavelength 266 nm (see Methods).
In the quantum realm, this classical system has its close counterpart, where each particle is represented by a localized wave-packet propagating without spreading along the resonant trajectory with a period $sT$ [29, 1]. In the quantum world, there is the tunneling process that can lead to the transfer of a particle from one wave-packet to neighboring wave-packets. Tunneling times become very long and thus negligible in the experiment if the wave-packets are sufficiently strongly localized. In the following, we consider ultra-cold atoms and assume negligible natural tunneling, but we can selectively introduce tunneling between any wave-packets in a controlled manner using Bragg scattering [40], as explained in Fig. 1b (see also Methods). In summary, this allows us to create a system of non-interacting bosons (or fermions) described by the tight-binding model,
$\hat{H}=\frac{1}{2}\sum_{i,j=1}^{s}J_{ij}\hat{a}_{i}^{\dagger}\hat{a}_{j},$ (1)
where $\hat{a}_{i}$ ($\hat{a}_{i}^{\dagger}$) represents the bosonic annihilation (creation) operator, responsible for annihilating (creating) a particle in the $i$-th wave-packet. All Bragg tunneling amplitudes $J_{ij}$ are subject to precise control and can be engineered at will at any moment during an experiment. This model describes a crystalline structure in time, or a time lattice. Unlike static states localized in a periodic spatial structure, the lattice sites here correspond to evolving localized wave-packets that form a periodic structure in time (see also Methods).
Initially, we assume non-interacting atoms. This scenario can be realized by means of a Feshbach resonance, i.e., an external magnetic field can be chosen so that the s-wave scattering length which describes the contact interactions between atoms is zero [44]. The interactions between atoms can then be selectively reinstated, even between atoms occupying different lattice sites in (1), i.e., with $i\neq j$. A similar array of laser beams as used for selective Bragg scattering and control of the tunneling amplitudes $J_{ij}$ can serve this purpose. These laser beams allow for a selective change in the internal state of the atoms by Raman transfer occupying given wave-packets [42]. This change leads to interaction between the atoms when their respective wave-packets intersect during the periodic evolution along the resonant trajectory because the scattering length is no longer zero, as depicted in Fig. 1c. Subsequent to the interaction, the atoms can be reverted to their initial internal state. This approach allows for precise control over interactions between atoms occupying any pair of wave-packets which are characterized by the coefficients $U_{ij}$ in the Hamiltonian (see Methods)
$\hat{H}=\frac{1}{2}\sum_{i,j=1}^{s}\left(J_{ij}\hat{a}_{i}^{\dagger}\hat{a}_{j}+U_{ij}\hat{a}_{i}^{\dagger}\hat{a}_{j}^{\dagger}\hat{a}_{j}\hat{a}_{i}\right).$ (2)
Hence, we have a system in which all tunneling amplitudes $J_{ij}$ and all interaction coefficients $U_{ij}$ can be controlled at will at any moment during an experiment.
The Hamiltonian (2) is derived for bosons but a similar universal Hamiltonian can be obtained for fermions. Fermions can be prepared in a time crystalline structure and their tunnelings and interactions can be controlled selectively.
@@
Figure 2: Temporal printed circuit board. (a): The $s$ wave-packets evolving along the resonant trajectory (Fig. 1a) can be treated as $s$ states of a lattice with $s$ sites. If we arrange the sites of the lattice in a 1D chain and using Bragg scattering (Fig. 1b) we realize tunneling between nearest neighbors, we will have a 1D crystalline structure with nearest neighbor hoppings. If we arrange the sites in a 2D lattice and induce tunneling between nearest neighbors, we will have a 2D crystalline structure. Similarly, we can realize 3D and higher-dimensional crystalline structures. (b): Let us consider $s$ periodically evolving wave-packets as states of a 2D lattice. Using a broad laser beam and having only two focused laser beams at our disposal (Fig. 1b), all nearest neighbor tunnelings of the 2D lattice can be realized. For example, a 2D sublattice can be coupled to another 2D sublattice via a 1D chain as shown in the figure (realized tunnelings are marked with short blue arrows). Control over the phases of tunneling amplitudes $J_{ij}$ allows for the generation of an artificial magnetic field in the system. With an array of focused laser beams available, any sites in the lattice can be connected via tunneling, as illustrated by the long blue arrow. More exotic geometries can also be realized, such as the Klein bottle, where the left and right edges of the 2D lattice are connected normally, but the top edge is twisted before connection to the bottom edge. Using selective Raman transfer (Fig. 1c), interactions between atoms occupying any pair of sites can be realized and controlled (as indicated by red arched arrows in the plot). During an experiment, the entire system can undergo any reconfiguration.
@@
With the universal time crystalline structure described by the Hamiltonian (2), it is possible to realize a broad range of quantum devices. Figure 2a presents a few examples. Physically, the states associated with the sites of the time crystalline lattice are localized wave-packets moving periodically in a 1D box (Fig. 1a). One can represent the crystalline structure as a row of sites but if the sites are connected via selective Bragg tunneling as depicted in the middle plot in Fig. 2a, then the structure can be represented as a 2D lattice with nearest neighbor tunnelings, and 2D condensed matter problems can be realized and investigated. For example, fermionic or bosonic atoms can experience a magnetic-like field if proper complex phases of the tunneling amplitudes $J_{ij}$ are realized by a proper choice of the relative phase of the laser beams in the Bragg scattering (see Methods). A few such 2D lattices can be created and connected via selective tunneling to form a 3D lattice (bottom plot in Fig. 2a), and this procedure can be continued to form a 4D lattice, and so on. It is also possible to realize exotic objects like a Klein bottle, which is hard to imagine in 3D space [45, 46]. That is, pairs of opposite edges of a 2D lattice can be connected via tunneling, but one of the edges is twisted before the connection.
@@
An entire time crystalline structure can be represented as sites on a 2D board, as shown in Fig. 2b. Such a board can be considered analogous to a printed circuit board in electronics. Indeed, any site on the board can be connected to any other site through selective tunneling, and an atom occupying any site can interact with an atom occupying any other site. All tunnelings and interactions can be individually turned on or off or modified because all parameters of the Hamiltonian (2) are fully controlled. Thus, in one part of the board, we can realize, for example, a 2D structure, while in other parts of the board, there are other 2D or higher-dimensional structures, or a structure with fully connected sites, or even a structure in the form of a Klein bottle. We can realize and control arbitrary connections between the structures and reconfigure the entire system at any moment during the experiment. Consequently, we can realize a broad range of quantum devices that can perform the operations we need.
@@
Figure 3: Quantum computer. (a): Initially, $s/2$ bosonic atoms are prepared in a Mott insulator phase in a 1D box potential in the presence of a static optical lattice. We assume there are $s/2$ potential wells in the box, with one atom in each well. After preparing the Mott insulator phase, the interactions between atoms are turned off by means of a Feshbach resonance [44]. The next step is to impart momentum to the atoms that satisfies the $s:1$ resonance condition with the frequency of the optical lattice oscillation, which we simultaneously switch on. Shortly after the previously described procedure, we observe $s/2$ wave-packets occupied by single atoms moving to the right in the figure and $s/2$ unoccupied wave-packets moving to the left. We assign one unoccupied wave-packet to each occupied wave-packet, forming the $|0\rangle$ and $|1\rangle$ states of qubits. In total, we have $s/2$ qubits. When two wave-packets corresponding to the $|0\rangle$ and $|1\rangle$ states of the same qubit meet during evolution along the resonant trajectory, a single-qubit gate can be performed using Bragg scattering. Control over the relative phase of the laser beams used in the Bragg scattering allows us to control whether the $\sigma_{x}$ or $\sigma_{y}$ operation is performed. High fidelity single-qubit operations are achieved by dividing them into several stages, meaning the entire single operation requires multiple encounters of wave-packets, and hence spanning several periods of the resonant trajectory (Fig. 1d). (b): When wave-packets corresponding to the states $|1\rangle_{i}$ and $|1\rangle_{j}$ of the $i$-th and $j$-th qubits meet during evolution, a controlled-Z gate can be realized. Atoms initially do not interact, but just before the wave-packets pass each other, we change the internal states of the atoms to states where they interact. We do this using Raman transfer with two beams. After the wave-packets pass each other, another Raman transfer restores the initial internal states of the atoms. If the interaction strength between atoms and the duration of interaction are appropriately chosen, after the wave-packets pass, the state $|1\rangle_{i}|1\rangle_{j}$ acquires a phase $e^{i\pi}$, and the controlled-Z gate is completed. Similarly to single-qubit operations, a high fidelity controlled-Z gate can be realized by dividing it into several encounters of the appropriate wave-packets (Fig. 1e-1f). (c): Quantum volume [47] calculated based on the B-gate decomposition of a generic SU(4) two-qubit operation [48]. In this plot we assume no error in the single-qubit operations. We plot a square of fidelilty of decomposition of a single B-gate using optimal points selected from the CZ-fidelity plot (Fig. 1e). For higher Raman transfer fidelity one can use a longer multi-cycle realization of the CZ-gates leading to quick increase in the quantum volume. The results correspond to the same parameters as in Figs. 1d-1f.

### v3

@@
Abstract
Time crystalline structures can be created in periodically driven systems. They are temporal lattices which can reveal different condensed matter behaviors ranging from Anderson localization in time to temporal analogues of many-body localization or topological insulators. However, the potential practical applications of time crystalline structures have yet to be explored. Here, we pave the way for time-tronics where temporal lattices are like printed circuit boards for realization of a broad range of quantum devices. The elements of these devices can correspond to structures of dimensions higher than three and can be arbitrarily connected and reconfigured at any moment. Moreover, our approach allows for the construction of a quantum computer, enabling quantum gate operations for all possible pairs of qubits. Our findings indicate that the limitations faced in building devices using conventional spatial crystals can be overcome by adopting crystalline structures in time.
I Introduction
@@
We first illustrate how resonantly driven ultra-cold atoms can create a time crystalline structure where the connection of arbitrary sites through the tunneling of atoms between sites or atom-atom interactions can be realized (Sec. II). As any connections between sites can be controlled, it is possible to realize a broad class of quantum devices, ranging from one-, two-, three- or higher-dimensional structures to more exotic objects like a Klein bottle, all of which can be connected in an arbitrary way, and the entire system can be reconfigured at any time during an experiment. Next, in Sec. III, we focus on a specific example, a quantum computer [39]. We demonstrate that a temporal printed circuit board can host qubits, where all single-qubit operations can be realized and a controlled-Z gate can be performed between all possible qubit pairs, meeting the conditions for a universal quantum computer. Section IV contains the summary and conclusions.
II Temporal printed circuit board
In this section, we describe how resonantly driven ultracold atoms can be used to construct a temporal printed circuit board — a device based on a temporal lattice, in which atomic tunneling between arbitrary lattice sites can be implemented, interactions between atoms occupying any chosen pairs of sites can be controlled, and the entire system can be dynamically reconfigured during the course of the experiment. We begin by introducing the concept of a temporal crystalline structure, then explain how tunneling and interactions can be controlled, and finally discuss the idea of the circuit board.
II.1 Time crystalline structure
@@
Figure 1: Controlling tunneling and interactions between resonantly driven quantum wave-packets. (a): Red ball illustrates a classical particle moving in a 1D box potential in the presence of an oscillating spatially periodic potential. If the period of particle motion is $s$ times longer than the period of the oscillating potential, $T=2\pi/\omega$, we have an $s:1$ resonance. In such a case, $s$ classical non-interacting particles can be positioned so that they follow one another along the resonant trajectory. In the quantum description, the $s:1$ resonance manifests itself as $s$ localized wave-packets, which evolve one after another along the resonant trajectory. Here, we assume that the wave-packets are sufficiently localized so that natural tunneling between neighboring wave-packets is negligible. (b): If during the encounter of the $i$-th and $j$-th wave-packets moving in opposite directions, one broad laser beam (labelled $\mathbf{k}_{2}$) and one narrow laser beam (labelled $\mathbf{k}_{1}$) focused on the meeting point of the wave-packets are briefly turned on and the Bragg scattering [40] condition is satisfied, we can realize atom tunneling between the wave-packets with an amplitude $J_{ij}$, the modulus of which depends on the beam parameters and the interaction time with the atom, and a phase depending on the relative phase between the beams. With a controllable array of focused laser beams available (e.g., using a digital micromirror device [41] as in the figure), which can be independently activated at different time points, one can control the tunneling amplitudes $J_{ij}$ between any pair of wave-packets. (c): Ultra-cold atoms can be prepared via a Feshbach resonance in states where they do not interact. In such a situation, we have the possibility of selectively controlling interactions between atoms occupying any pair of wave-packets. If just before the encounter of atoms occupying two wave-packets, we perform a Raman transfer [42] (using broad and focused laser beams) from the initial internal states of the atoms to states where the atoms interact, then during the passage of the wave-packets, interaction between atoms will occur. After the wave-packets pass each other, we perform a Raman transfer again, but back to non-interacting internal states of the atoms. With a digital micromirror device available, interaction between atoms occupying any pair of wave-packets can be realized.
Consider a single atom confined within a one-dimensional (1D) box potential periodically perturbed by an oscillating optical lattice potential [Fig. 1(a)]. Using units of length, energy, and time, denoted by $1/k$, $\hbar^{2}k^{2}/m$, and $m/\hbar k^{2}$, respectively, where $k$ is the wave number of the laser beam creating the oscillating optical lattice potential, and $m$ is the mass of the atom, the system Hamiltonian is given by
Before considering the quantum resonant behavior of the atom, let us commence with a classical description. We introduce a new pair of momentum and position variables, known as action-angle variables [26], $I=s|p|/2$ and $|\theta|=2z/s$ where $\pi\leq\theta<\pi$. Switching to the moving frame, $\Theta=\theta-\omega t/s$, and neglecting rapidly oscillating terms, the effective Hamiltonian becomes $H_{\rm eff}=P^{2}/2+(\lambda s^{2}/32)\cos(s\Theta)$, where $P=I-\omega/s$ [26, 27]. For $s\gg 1$, $H_{\rm eff}$ describes an atom moving in a spatially periodic potential. In the quantum description, this results in solid-state-like behavior, which is manifested as the formation of energy bands and eigenstates in the form of Bloch waves [27, 28].
In the following, we confine our focus to the lowest energy band of the quantum version of the Hamiltonian $H_{\rm eff}$. The corresponding Hilbert subspace can be expanded in the Wannier state basis, $w_{i}(\Theta)$, where $w_{i}(\Theta)$ is localized at the $i$-th site of the potential in $H_{\rm eff}$ [27, 29]. In the laboratory frame, these are localized wave-packets, $w_{i}(z,t)$, evolving along the periodic resonant trajectory with the resonant momentum $p$, i.e., with the period $sT$, where $T=2\pi/\omega$ [Fig. 1(a)]. In this basis, the Hamiltonian of many non-interacting bosonic atoms takes the form
$\hat{H}=\frac{J}{2}\sum_{i=1}^{s}\hat{a}_{i+1}^{\dagger}\hat{a}_{i}+\rm{h.c.},$ (2)
where $\hat{a}_{i}$ ($\hat{a}_{i}^{\dagger}$) represents the bosonic annihilation (creation) operator, responsible for annihilating (creating) a particle in the $i$-th wave-packet and $J=2\int_{0}^{L}dz,w_{i+1}^{*}(H_{0}-i\partial_{t})w_{i}$ represents the tunneling amplitude of atoms between the neighboring wave-packets on the resonant trajectory [27, 43]. We set the oscillation amplitude $\lambda$ such that these natural tunnelings are negligible ($J\approx 0$), as our intention is to selectively introduce tunneling in a controlled manner.
II.2 Control of the tunneling amplitudes
To selectively control atom transfer between different wave-packets, we can utilize Bragg scattering [40]. Let us consider two laser beams characterized by wave vectors $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ that we switch on for a short period of time $\tau$. The first beam is focused at the location where two wave-packets, $w_{i}(z,t)$ and $w_{j}(z,t)$, meet during evolution [Fig. 1(b)]. The waist of the beam is larger than the width of the wave-packets but smaller than the distance between potential minima in (1). The second beam is broad and covers the entire 1D box. If the Bragg condition is satisfied, $\mathbf{k}_{1}-\mathbf{k}_{2}=2p\mathbf{e}_{z}$, an atom occupying the wave-packet $w_{j}$, moving with an average momentum $p\mathbf{e}_{z}$, is transferred to the other wave-packet $w_{i}$, which moves with an average momentum $p\mathbf{e}_{z}$, and the reverse transfer is equally probable. These processes are described by an additional term in the Hamiltonian [40]
$H_{B}=\lambda_{\rm Bragg}(t)\cos^{2}\left[\frac{(\mathbf{k}_{1}-\mathbf{k}_{2})\cdot\mathbf{e}_{z}z+\phi}{2}\right]\frac{e^{-\frac{(z-z_{0})^{2}}{W^{2}}}}{\sqrt{4\pi W^{2}}},$ (3)
where $z_{0}$ and $W$ are the location and waist of the narrow laser beam, respectively. We choose $W=\pi/(2\sqrt{2})$. $\lambda_{\rm Bragg}$ is determined by the intensities of the beams and $\phi$ is the relative phase of the beams. The beams are activated only for a short period of time $\tau$ around the moment when two relevant wave-packets pass each other. The resulting tunneling amplitude of the atom between the wave-packets is described by
$J_{ij}(t)=2\int_{0}^{L}dz\;w_{i}^{*}(z,t)\;H_{B}(z,t)\;w_{j}(z,t).$ (4)
The magnitude of $J_{ij}$ is controlled by the parameters of the beams and the duration of their interaction with the atom (in the following we choose $\tau=0.24T$, where $T=2\pi/\omega$, around central moments of wave-packet encounters), while the phase is controlled by the relative phase $\phi$. By switching on the beams during meetings of different wave-packets and having multiple focused beams available [Fig. 1(b)], we have the ability to control the transfer of atoms between all wave-packets, i.e., we can realize the Hamiltonian
$\hat{H}=\frac{1}{2}\sum_{i,j=1}^{s}J_{ij}(t)\;\hat{a}_{i}^{\dagger}\hat{a}_{j}.$ (5)
II.3 Control of the interaction strengths
Interactions between ultra-cold atoms are effectively short-range contact interactions described by a single parameter, the s-wave scattering length [44]. In the Hilbert subspace spanned by the wave-packets, $w_{i}(z,t)$, evolving along the periodic resonant trajectory, the Hamiltonian of interacting atoms reads [27, 43]
$\hat{H}(t)=\frac{1}{2}\sum_{i,j=1}^{s}\left(J_{ij}(t)\;\hat{a}_{i}^{\dagger}\hat{a}_{j}+U_{ij}(t)\;\hat{a}_{i}^{\dagger}\hat{a}_{j}^{\dagger}\hat{a}_{j}\hat{a}_{i}\right),$ (6)
with the interaction coefficients,
$U_{ij}(t)=g_{0}(t)\int_{0}^{L}dz\;|w_{i}(z,t)|^{2}\;|w_{j}(z,t)|^{2},$ (7)
where $g_{0}(t)=2m\omega_{\perp}a(t)/\hbar k$ and $a(t)$ is the scattering length which can be changed in the experiment. Choosing properly an external magnetic field in the vicinity of a Feshbach resonance [44], we can set the scattering length $a=0$. If we change the internal state of the atoms that occupy two wave-packets which are about to meet during the evolution along the periodic trajectory, they will interact because the scattering length $a$ is no longer zero. The change of internal state can be performed by a focused laser beam which together with another broad beam realizes a Raman transition between the internal states [42], see Fig. 1(c). Immediately after the wave-packets pass each other, the laser beams can be applied once again to restore the initial internal states of the atoms. The entire procedure leads to a time-dependent $g_{0}(t)$ which is non-zero only during the passing of the wave-packets. With an array of focused laser beams one can control and engineer all interaction coefficients $U_{ij}$ in the Hamiltonian (6).
II.4 The idea of a temporal printed circuit board
We have demonstrated that ultra-cold atoms confined in a box potential and driven by a periodically oscillating optical lattice can be described by the Hamiltonian (6). In this framework, it is possible to engineer and precisely control tunneling amplitudes, $J_{ij}$, between arbitrary sites via Bragg scattering, as well as tune the interaction strengths, $U_{ij}$, between atoms occupying any selected pair of sites and everything can be modified at any moment during the experiment. The Hamiltonian (6) is derived for bosons, but a similar universal Hamiltonian can be obtained for fermions.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ II.4 The idea of a temporal printed circuit board
Figure 2: Temporal printed circuit board. (a): The $s$ wave-packets evolving along the resonant trajectory [Fig. 1(a)] can be treated as $s$ states of a lattice with $s$ sites. If we arrange the sites of the lattice in a 1D chain and using Bragg scattering [Fig. 1(b)] we realize tunneling between nearest neighbors, we will have a 1D crystalline structure with nearest neighbor hoppings. If we arrange the sites in a 2D lattice and induce tunneling between nearest neighbors, we will have a 2D crystalline structure. Similarly, we can realize 3D and higher-dimensional crystalline structures. (b): Let us consider $s$ periodically evolving wave-packets as states of a 2D lattice. Using a broad laser beam and having only two focused laser beams at our disposal [Fig. 1(b)], all nearest neighbor tunnelings of the 2D lattice can be realized. For example, a 2D sublattice can be coupled to another 2D sublattice via a 1D chain as shown in the figure (realized tunnelings are marked with short blue arrows). Control over the phases of tunneling amplitudes $J_{ij}$ allows for the generation of an artificial magnetic field in the system. With an array of focused laser beams available, any sites in the lattice can be connected via tunneling, as illustrated by the long blue arrow. More exotic geometries can also be realized, such as the Klein bottle, where the left and right edges of the 2D lattice are connected normally, but the top edge is twisted before connection to the bottom edge. Using selective Raman transfer [Fig. 1(c)], interactions between atoms occupying any pair of sites can be realized and controlled (as indicated by red arched arrows in the plot). (c): Example building block of two connected $\operatorname{SU}(2)$ systems. The number of bosonic atoms decides the representation in each system $i$ — if there is one atom, we deal with the spin-1/2 system. Blue vertical arrows correspond to tunneling (ladder operators), the (anti-) ferromagnetic interactions can be supplied by programmed interactions depicted by horizontal or diagonal lines. (d): Example of 1D Bose-Hubbard system of 5 sites (labelled 1 to 5 in the plot) connected to two engineered reservoir systems. The reservoirs represented by all-to-all connected systems allow for extremely fast mixing and their spectral density can be designed. During an experiment, the parameters can undergo arbitrary reconfigurations. Here, the purely dephasing interactions $U(t)$ with one reservoir and particle exchange with the other reservoir $J(t)$ can be, e.g., periodically switched during the experiment allowing for study of quantum thermodynamics in a fully closed system.
@@
With the universal time crystalline structure described by the Hamiltonian (6), it is possible to realize a broad range of quantum devices. Figure 2 presents a few examples. Physically, the states associated with the sites of the time crystalline lattice are localized wave-packets moving periodically in a 1D box [Fig. 1(a)]. One can represent the crystalline structure as a row of sites but if the sites are connected via selective Bragg tunneling as depicted in the middle plot in Fig. 2(a), then the structure can be represented as a 2D lattice with nearest neighbor tunnelings, and 2D condensed matter problems can be realized and investigated. For example, bosonic or fermionic atoms can experience a magnetic-like field if proper complex phases of the tunneling amplitudes $J_{ij}$ are realized by a proper choice of the relative phase, $\phi$, of the laser beams in the Bragg scattering, see (3). A few such 2D lattices can be created and connected via selective tunneling to form a 3D lattice [bottom plot in Fig. 2(a)], and this procedure can be continued to form a 4D lattice, and so on. It is also possible to realize exotic objects like a Klein bottle, which is hard to imagine in 3D space [45, 46]. That is, pairs of opposite edges of a 2D lattice can be connected via tunneling, but one of the edges is twisted before the connection.
@@
An entire time crystalline structure can be represented as sites on a 2D board, as shown in Fig. 2(b). Such a board can be considered analogous to a printed circuit board in electronics. Indeed, any site on the board can be connected to any other site through selective tunneling, and an atom occupying any site can interact with an atom occupying any other site. All tunnelings and interactions can be individually turned on or off or modified because all parameters of the Hamiltonian (6) are fully controlled. Thus, in one part of the board, we can realize, for example, a 2D structure, while in other parts of the board, there are other 2D or higher-dimensional structures, or a structure with fully connected sites, or even a structure in the form of a Klein bottle. We can realize and control arbitrary connections between the structures and reconfigure the entire system at any moment during the experiment. Consequently, we can realize a broad range of quantum devices that can perform the operations we need — examples are presented in Figs. 2(c)-2(d).
Since the coupling coefficients $J_{ij}(t)$ and $U_{ij}(t)$ are activated at certain moments of time, the entire Hamiltonian (6) is time-dependent. However, if we apply a sequence of the Bragg and Raman pulses in the same way every period of the resonant trajectory and if
$\frac{1}{|J_{ij}|}\gg sT,\quad\frac{1}{|U_{ij}|N/s}\gg sT,$ (8)
we may derive the effective time-independent Hamiltonian. That is, if the timescales associated with atomic tunneling and interactions between atoms are much longer than the period of the resonant trajectory, it is justified to perform a time average of the original Hamiltonian, yielding the effective Hamiltonian
$\hat{H}_{\rm eff}=\frac{1}{2}\sum_{i,j=1}^{s}\left(\bar{J}_{ij}\;\hat{a}_{i}^{\dagger}\hat{a}_{j}+\bar{U}_{ij}\;\hat{a}_{i}^{\dagger}\hat{a}_{j}^{\dagger}\hat{a}_{j}\hat{a}_{i}\right),$ (9)
where
$\bar{J}_{ij}$ $=$ $\frac{1}{sT}\int_{0}^{sT}dt\;J_{ij}(t),$ (10)
$\bar{U}_{ij}$ $=$ $\frac{1}{sT}\int_{0}^{sT}dt\;U_{ij}(t).$ (11)
The circuit board is now described by a time-independent Hamiltonian $\hat{H}_{\rm eff}$, yet at any moment during the experiment we can fully reprogram the sequence of Bragg and Raman pulses, thereby driving the system to evolve within an entirely different effective structure. This enables the realization of a sequence of different quantum operations on a system of ultra-cold atoms. The initial operations may be used to prepare an initial state of the many-body system, while subsequent ones can perform the desired computational tasks.
We describe the realization of the temporal printed circuit board in ultra-cold atoms moving between the walls of a box potential. This potential offers significant advantages, as the laser beam parameters required for Bragg scattering can remain identical regardless of where the evolving wave-packets meet. Box potentials are implemented in ultra-cold atomic gas laboratories and are used in experimental studies of many-body physics [47]. However, we emphasize that temporal printed circuit boards can also be realized in other trapping potentials and also for atoms bouncing on an atomic mirror [48]. Any trapping potential that is not harmonic allows for stable resonant atomic motion. Similar to the case of the box potential, the same tools (Bragg scattering and Raman transitions) can be used to implement the temporal printed circuit board. The only difference is that atoms evolving along a resonant trajectory will have different momenta at different positions due to variations in potential energy. Consequently, the parameters of Bragg pulses must be adjusted at different points where evolving wave-packets meet.
@@
Figure 3: Quantum computer. (a): Initially, $s/2$ bosonic atoms are prepared in a Mott insulator phase in a 1D box potential in the presence of a static optical lattice. We assume there are $s/2$ potential wells in the box, with one atom in each well. After preparing the Mott insulator phase, the interactions between atoms are turned off by means of a Feshbach resonance [44]. The next step is to impart momentum to the atoms that satisfies the $s:1$ resonance condition with the frequency of the optical lattice oscillation, which we simultaneously switch on. Shortly after the previously described procedure, we observe $s/2$ wave-packets occupied by single atoms moving to the right in the figure and $s/2$ unoccupied wave-packets moving to the left. We assign one unoccupied wave-packet to each occupied wave-packet, forming the $|0\rangle$ and $|1\rangle$ states of qubits. In total, we have $s/2$ qubits. When two wave-packets corresponding to the $|0\rangle$ and $|1\rangle$ states of the same qubit meet during evolution along the resonant trajectory, a single-qubit gate can be performed using Bragg scattering. Control over the relative phase of the laser beams used in the Bragg scattering allows us to control whether the $\sigma_{x}$ or $\sigma_{y}$ operation is performed. High fidelity single-qubit operations are achieved by dividing them into several stages, meaning the entire single operation requires multiple encounters of wave-packets, and hence spanning several periods of the resonant trajectory, see Fig. 4(a). (b): When wave-packets corresponding to the states $|1\rangle_{i}$ and $|1\rangle_{j}$ of the $i$-th and $j$-th qubits meet during evolution, a controlled-Z gate can be realized. Atoms initially do not interact, but just before the wave-packets pass each other, we change the internal states of the atoms to states where they interact. We do this using Raman transfer with two beams. After the wave-packets pass each other, another Raman transfer restores the initial internal states of the atoms. If the interaction strength between atoms and the duration of interaction are appropriately chosen, after the wave-packets pass, the state $|1\rangle_{i}|1\rangle_{j}$ acquires a phase $e^{i\pi}$, and the controlled-Z gate is completed. Similarly to single-qubit operations, a high fidelity controlled-Z gate can be realized by dividing it into several encounters of the appropriate wave-packets, see Figs. 4(b)-4(c).
III Quantum computer
As a compelling demonstration of the versatility of our temporal printed circuit board, we explore its application to the implementation of a quantum computer. Although not the central focus of this work, this represents one of the most experimentally challenging architectures, making it an ideal and stringent benchmark for showcasing both the power and the fundamental limits of our approach.
III.1 The idea of realizing a quantum computer